OpenStudy (crashonce):

In order to solve the equation log x^2+2x=0, what line must be added to the graph of y=logx-1

OpenStudy (crashonce):

A) y=-x B) y=-x-1 C) y=-x+1 D) y=2x-1 E) y=2x+1

OpenStudy (crashonce):

@hartnn

OpenStudy (crashonce):

@ganeshie8 @mathmath333

hartnn (hartnn):

log (x^2+2x) - 0 ?

OpenStudy (crashonce):

not sure wat u mean

hartnn (hartnn):

lol, whats the question ?

hartnn (hartnn):

log (x^2+2x) = 0 ?

OpenStudy (crashonce):

no logx^2 + 2x outside

hartnn (hartnn):

ok, so \(\log (x^2)+2x =0 \)

hartnn (hartnn):

log x^n = n log x

OpenStudy (michele_laino):

I think that: from \[\log (x ^{2})+2x=0,\] I get: \[\log (x)+x=0\] and: \[\log(x)-1+1+x=0\] from which: \[y+1+x=0\]

OpenStudy (michele_laino):

nevertheless the there is not the option 1+x

hartnn (hartnn):

log x + x = 0 subtract log x -1 =0 you get x - (-1) = 0 x+1 = 0 yeah, i get y=x+1 too or y = -x-1 will also do

OpenStudy (crashonce):

thanks both

OpenStudy (michele_laino):

thank you! @CrashOnce

hartnn (hartnn):

welcome ^_^